OpenStudy (anonymous):

Solve this system of equations. Please write answer as an ordered pair. y=x 2x+5y=28

OpenStudy (kagıtucak):

y=x y-x=0 (1) 2x+5y=28 (2) Multiply (1) with 2 and sum up (1)*2 and (2) At this moment, x will dissapear and we get 7y=28 y=4 Put it into (1) 4-x=0 x=4 Thus our answer is (4,4)

OpenStudy (anonymous):

Thank you so much!

OpenStudy (kagıtucak):

Your're welcome!